how do you simply ln(1/e^2)
Answer by ikleyn(52802)

.
ln(1/e^2) = ln(1) - ln(e^2) = 0 - 2*ln(e) = 0 - 2*1 = 0 - 2 = -2.


On logarithms and their properties, see the lessons
    - WHAT IS the logarithm
    - Properties of the logarithm
    - Change of Base Formula for logarithms
    - Solving logarithmic equations
    - Using logarithms to solve real world problems
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Logarithms".